OpenStudy (anonymous):

find all the real cube roots of -0.000125

OpenStudy (anonymous):

Medal and fan :3

jhonyy9 (jhonyy9):

do you know cuberoot(-1) = ?

OpenStudy (anonymous):

i dont

jhonyy9 (jhonyy9):

and 0,000125 you can rewrite it = 125*10^(-6) --- yes ?

OpenStudy (anonymous):

mhmmm

jhonyy9 (jhonyy9):

cuberoot x = ? using exponential form ?

jhonyy9 (jhonyy9):

hope you know that 125=5*5*5 than how many will be cuberoot 125 = ?

jhonyy9 (jhonyy9):

so because -1*(-1)*(-1)=-1 from this result that cuberoot(-1)=-1 yes ?

OpenStudy (anonymous):

yes i understand

jhonyy9 (jhonyy9):

so now you have got the result for cuberoot(-125)= ?

OpenStudy (anonymous):

thank you for helping me out. i was completely lost but it makes a little more sense now

jhonyy9 (jhonyy9):

so this mean that you know the right answer how many will be ?

OpenStudy (anonymous):

it equals -5x-5x-5

jhonyy9 (jhonyy9):

but from this result just -125 and your exercise is - 0,000125 inside cuberoot - or not ?

OpenStudy (anonymous):

i got 0.0025 and -0.0025 as the answers

jhonyy9 (jhonyy9):

you have got it now this result so than you understand it sure right ,yes ?

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

ok. than was my pleasure good luck bye

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

welcome bye